One positive number Is 5 times another number. The difference between the two numbers is 304. Find the numbers
Answer by Shana-D77(132)

"One positive number Is 5 times another number."
Ok, so let the numbers be x and y.
I'll say that x is the bigger number:
x = 5y
"The difference between the two numbers is 304. Find the numbers."
x - y = 304
Because x = 5y is solved for x, let's solve the second equation for x, too.
x = 304 + y (aded y to both sides)
Because both equations are now solved for x (both =x), they're equal to each other by the transitive property. So,
5y = 304 + y
4y = 304 (subtracted y from both sides)
y = 76
Because x = 5y, x = 5(76) = 380
(380, 76)